### v1

# SE4Lip: Speech-Lip Encoder for Talking Head Synthesis to Solve Phoneme-Viseme Alignment Ambiguity

Abstract
Speech-driven talking head synthesis tasks commonly use general acoustic features (such as HuBERT and DeepSpeech) as guided speech features. However, we discovered that these features suffer from phoneme-viseme alignment ambiguity, which refers to the uncertainty and imprecision in matching phonemes (speech) with visemes (lip). To address this issue, we propose the Speech Encoder for Lip (SE4Lip) to encode lip features from speech directly, aligning speech and lip features in the joint embedding space by a cross-modal alignment framework. The STFT spectrogram with the GRU-based model is designed in SE4Lip to preserve the fine-grained speech features. Experimental results show that SE4Lip achieves state-of-the-art performance in both NeRF and 3DGS rendering models. Its lip sync accuracy improves by 13.7% and 14.2% compared to the best baseline and produces results close to the ground truth videos.
Figure 1: The comparison of the talking head synthesis pipeline when using acoustics features and SE4Lip. The core of SE4Lip is to solve phoneme-viseme alignment ambiguity, which refers to the uncertainty and imprecision in matching phonemes (speech) with visemes (lip). SE4Lip aligns lip features with speech using a cross-modal alignment framework. Thus, SE4Lip can significantly improve the quality of the synthesized videos.
1 Introduction
Talking head synthesis has attracted widespread applications in video conferences [41], film production [18], psychology [8], and other fields. There is an expectation to generate dynamic, realistic, and stable synthetic videos, especially regarding lip movements. The pipeline of talking head synthesis is illustrated in Fig. 1. Speech and visual features are used as conditional inputs to a rendering model, which ultimately synthesizes the video. Since this task is driven by speech, the quality of the synthesized video heavily depends on the quality of speech features.
[…]
This paper proposes SE4Lip to solve the issue of phoneme-viseme alignment ambiguity. SE4Lip is based on the idea of contrastive learning and employs a cross-modal alignment framework to address this issue. SE4Lip models the speech signal using a combination of the STFT spectrogram and the GRU-based model instead of the traditional approach that combines the Mel spectrogram and the CNN-based model [37]. This is because, compared to the STFT spectrogram, the Mel spectrogram compresses the feature space in the frequency domain, leading to the loss of fine-grained frequency information. Additionally, the GRU is more effective in capturing temporal variations, thereby enhancing the ability of speech features to represent lip movements. Inspired by Wav2Lip [30] and SyncNet [7], SE4Lip uses a CNN-based model to extract lip features. The contrastive loss function forces SE4Lip to focus on the causal relationship between phonemes and lip movements rather than linguistic representations. As shown in Fig. 1 with the orange arrows, SE4Lip strengthens the alignment capability of the rendering model, allowing it to synthesize accurate and clear lips. Experimental results show that videos synthesized using SE4Lip outperform those synthesized using acoustics features in different rendering models. We summarize our contributions as follows.
  1. 1.
    Speech-Lip Encoder: To address the phoneme-viseme alignment ambiguity issue, we train a speech encoder specially for talking head synthesis through a cross-modal alignment framework. This framework directly establishes the alignment between speech and lip features in a shared feature space, effectively avoiding the issue of phoneme-viseme alignment ambiguity.
  2. 2.
    Fine-Grained Speech Signal Processing: To preserve fine-grained information in speech features, we propose combining the STFT spectrogram and the GRU model for speech feature extraction. The STFT spectrogram retains more frequency details, while the GRU effectively captures the temporal relationships. This combination of STFT and GRU significantly improves the accuracy and clarity of the lip shapes.
  3. 3.
    Excellent Performance: Compared to four acoustics features, we achieve state-of-the-art results on both NeRF and 3DGS rendering models. Our method achieves a 13.7% improvement in lip sync error confidence and a 14.2% improvement in lip sync error distance compared to the best baseline. The results are also closely aligned with ground truth videos. Furthermore, the results of the ablation experiments validate the effectiveness of the STFT spectrogram and the GRU-based model.
2 Related Work
2.1 Acoustics Feature
Acoustics features aim to extract linguistic representations from speech signals for subsequent tasks such as speech recognition and speaker identification. Representative works include DeepSpeech [1], HuBERT [15], Wav2Vec 2.0 [3] and Whisper [31].
Classification Manner of Pronunciation Examples
[…]
Affricate Complete oral closure and slow release of the lung air // & //
Table 1: The list of some phonemes (using consonants as an example) related to phoneme-viseme alignment ambiguity. “&” indicates that the lip shapes of the two phonemes are similar.
  1. 1.
    DeepSpeech [1]. DeepSpeech is a speech feature extraction model based on an end-to-end recurrent neural network. This model first combines the CTC loss function with a multi-layer bidirectional LSTM, directly mapping the Mel spectrogram to character sequences.
  2. 2.
    HuBERT [15]. HuBERT is a self-supervised learning framework. HuBERT obtains latent acoustic units through downsampling via convolutional layers and uses a BERT-style mask strategy for unit prediction.
  3. 3.
    Wav2Vec 2.0 [3]. Wav2Vec 2.0 is a speech feature extraction model based on a convolutional-transformer hybrid architecture. Wav2Vec 2.0 converts raw waveforms into latent vectors using a quantization module combined with contrastive loss for pretraining.
  4. 4.
    Whisper [31]. Whisper employs an encoder-decoder structure to jointly train speech recognition and language identification tasks. Mel spectrograms in Whisper are processed through transformer encoder blocks, which include convolutional layers with GELU [14] activations.
Although these features perform excellently in downstream tasks such as speech recognition and speaker identification, they are not designed with the temporal and frequency detail requirements of the talking head synthesis task in mind. They are designed to extract linguistic representations from the speech signal but cannot enhance the rendering model’s alignment capability.
Figure 2: The lip shape diagram of some phonemes related to phoneme-viseme alignment ambiguity.
2.2 Talking Head Synthesis
In recent years, due to its applications in digital humans, virtual avatars, and video conferencing, talking head synthesis [27, 36, 45, 9, 40, 11, 22, 29, 43, 2, 23, 16, 34, 13, 26, 4], especially the real-time talking head synthesis [38, 11, 21, 29, 43, 42, 20, 24, 25, 5, 12, 2, 23, 19], has attracted significant attention. The main rendering models currently used are based on Neural Radiance Fields (NeRF) [28] and 3D Gaussian Splatting (3DGS) [17]. NeRF achieves high-fidelity scene rendering by constructing an implicit continuous volume scene representation and modeling the color and density distribution of light propagation using multi-layer perceptions. 3D Gaussian Splatting, on the other hand, parameterizes the scene geometry and appearance using dynamic Gaussian point clouds and realizes real-time dynamic rendering through differentiable rasterization. In some representative works, AD-NeRF [11], ER-NeRF [22] and TalkingGaussian [23] use DeepSpeech as the speech feature; GeneFace [43] uses HuBERT; GaussianSpeech [2] uses Wav2Vec 2.0; and Salehi et al. [33] use Whisper.
Acoustic features are optimized for phoneme classification, focusing on the acoustic distinction between phonemes rather than the alignment relationship between phonemes and visemes. In the pipeline of traditional methods, phoneme-viseme alignment relies solely on the rendering model. However, existing rendering models often align phonemes and visemes through simple feature addition or attention mechanisms, which can not avoid the phoneme-viseme alignment ambiguity.
3 Motivation
Figure 3: The vocal tract anatomy diagram. The lips are only one of the organs that affect pronunciation.
3.1 Phoneme-Viseme Alignment Ambiguity
Phoneme-viseme alignment ambiguity refers to the phenomenon that different phonemes can correspond to a similar lip shape. In Tab. 1, we list some phonemes related to phoneme-viseme alignment ambiguity along with their corresponding pronunciation manner. Additionally, we present the specific lip shapes associated with the phoneme-viseme alignment ambiguity phenomenon in Fig. 2. To explain the origin of this phenomenon in detail, we also present a vocal tract anatomical diagram in Fig. 3 from Ramoo [32] to illustrate the organs that affect pronunciation. The larynx or vocal cords are the basis of pronunciation, while the lips and tongue are the articulatory organs. Although the lips are only a part of the speech production system, they are the most intuitive visual feature, which leads to the phoneme-viseme alignment ambiguity issue. For example, as shown in Fig. 4, the lip shapes corresponding to phonemes /d/ and /t/ are similar. However, /d/ is a voiced consonant with strong vocal cord vibrations during pronunciation, while /t/ is a consonant with weak vocal cord vibrations. Despite the significant acoustic differences between /d/ and /t/, the speech features of /d/ and /t/ should be aligned to similar visual features in the cross-modal alignment.
Figure 4: Taking the phonemes /d/ and /t/ as examples to illustrate the phoneme-viseme alignment ambiguity. In the visualization of HuBERT [15] features, there is a noticeable difference between /d/ and /t/. In the visualization of SE4Lip features, there is no significant distinction between /d/ and /t/ because they share similar lip shapes, indicating that SE4Lip differentiates phonemes based on lip shapes rather than acoustic features.
3.2 Drawback of Using Acoustics Features
Acoustics-driven talking head synthesis suffers from the issue of phoneme-viseme alignment ambiguity. The root cause is that acoustics features do not provide additional alignment support in the rendering model, forcing the model to rely solely on its weak alignment, which is achieved through simple feature addition or attention mechanisms. In addition, the goals of acoustics features and talking head synthesis are conflicting. The former aims to extract highly discriminative linguistic representations from speech signals, while the latter seeks to align phonemes with lip shapes. As shown in the scatter plot of Fig. 4, we extracted HuBERT [15] features for /d/ and /t/ and visualized them using PCA [35] and t-SNE [39] methods. It is evident that due to the difference in linguistic representations, HuBERT strongly distinguishes between the two phonemes. However, the talking head synthesis task requires alignment between the phoneme and visual features (particularly the lip shape). Furthermore, acoustics features commonly use the Mel spectrogram as input, leading to the loss of fine-grained frequency information, which is critical for synthesizing lip movements.
Figure 5: Overview of the SE4Lip framework. For a given speech-lip pair, SE4Lip generates embeddings through the corresponding encoder. In the joint embedding space, SE4Lip pulls positive pairs closer and pushes negative pairs farther apart.
In general, using acoustics features to drive talking head synthesis results in the following drawbacks: 1) Misalignment of Task Objectives. The purpose of acoustic features is to align speech with linguistic representations, which fundamentally contradicts the requirement of aligning speech with lips in talking head synthesis. This contradiction exacerbates the phoneme-viseme alignment ambiguity, reducing the synthesized video’s lip shape accuracy. 2) Lack of Dynamic Representation. The temporal resolution of acoustics features (typically 25ms frame length) struggles to capture the instantaneous changes in lip movements. For example, the lip closure-opening process for the plosive sound /p/ lasts around 80–120ms. However, the dynamic features of this movement are dropped by the frequency compression of the Mel spectrogram, leading to a blurry lip shape in the synthesis.
3.3 Contrastive Learning of Speech-Lip
To enhance the alignment capability of the rendering model and address the issue of phoneme-viseme alignment ambiguity, we propose using cross-modal speech features as input for the rendering model. To train such a speech encoder, we use contrastive learning and directly establish the alignment between speech and lip features in a joint embedding space. Our approach forces the speech encoder to focus on the causal relationship between phonemes and visemes rather than speech content or speaker identity. Additionally, we preserve fine-grained speech feature information through detailed processing in both the frequency and time domains. Specifically, we propose using the STFT spectrogram instead of the Mel spectrogram as input to the speech encoder to avoid frequency compression. We also introduce a temporal model to improve the temporal resolution of speech features.
4 Methodology
4.1 Overview
The framework of SE4Lip is shown in Fig. 5. SE4Lip consists of a speech feature extractor and a lip feature extractor. Specifically, in the preprocessing stage, SE4Lip applies STFT transformation to the raw speech to obtain the spectrogram and extract the lip from the raw image. The speech feature extractor employs a GRU-based model [6], while the lip feature extractor uses a CNN-based model. Finally, SE4Lip uses the contrastive loss function to minimize the distance between positive pairs and maximize the distance between negative pairs. With this design, SE4Lip can efficiently capture the dynamic relationship between speech and lip movements. It is worth noting that SE4Lip is trained using continuous speech and image data within a fixed window size, which helps the model capture the dynamic changes in lip movements.
Layer Type Input Chanel Output Chanel Kernel Size Stride Padding
Conv2d 15 32 7×7 1 3
Conv2d 32 64 5×5 (1,2) 1
ResidualBlock 64 64 3×3 1 1
ResidualBlock 64 64 3×3 1 1
Conv2d 64 128 3×3 2 1
ResidualBlock 128 128 3×3 1 1
ResidualBlock 128 128 3×3 1 1
Conv2d 128 256 3×3 2 1
ResidualBlock 256 256 3×3 1 1
Conv2d 256 512 3×3 2 1
ResidualBlock 512 512 3×3 1 1
Conv2d 512 512 3×3 2 1
Conv2d 512 512 3×3 (4,1) 0
Conv2d 512 512 1×1 1 0
Table 2: The network architecture of the lip feature extractor.
4.2 Speech Feature Extractor
SE4Lip generates the spectrogram using the Short-Time Fourier Transform (STFT). The Mel spectrogram compresses mid-to-high frequency features to simulate the human perception of frequency. Compared to the Mel spectrogram, the frequency distribution of STFT is linear, which preserves more fine-grained frequency information and captures detailed information related to lips. Traditional methods often use the Mel spectrogram, which loses some important frequency information during frequency compression. However, this information is vital for mapping the relationship between speech and lips.
PSNR↑(vs. HuBERT) LPIPS↓(vs. HuBERT) LMD↓(vs. HuBERT) LSE-C↑(vs. HuBERT) LSE-D↓(vs. HuBERT)
Ground Truth N/A 0 0 8.8302 (+15.3%) 6.0570 (-15.2%)
HuBERT [15] 32.0108 (+0%) 0.0427 (+0%) 2.9616 (+0%) 7.6599 (+0%) 7.1402 (+0%)
DeepSpeech [1] 32.0338 (+0.07%) 0.0417 (-2.34%) 3.0321 (+2.38%) 7.4268 (-3.04%) 7.5418 (+5.63%)
NeRF [29] Wav2vec 2.0 [3] 31.6359 (-1.17%) 0.0422 (-1.17%) 3.7204 (+25.6%) 3.3782 (-55.9%) 10.200 (+42.8%)
Whisper [31] 31.2311 (-2.44%) 0.0436 (+2.11%) 3.6727 (+24.0%) 3.3416 (-56.4%) 10.202 (+42.8%)
SE4Lip(Ours) 32.2301 (+0.68%) 0.0399 (-6.56%) 2.8725 (-3.01%) 8.7098 (+13.7%) 6.1255 (-14.2%)
Ground Truth N/A 0 0 8.8302 (+74.3%) 6.0570 (-34.2%)
HuBERT [15] 30.9339 (+0%) 0.0411 (+0%) 2.9634 (+0%) 5.0660 (+0%) 9.2126 (+0%)
DeepSpeech [1] 30.8336 (-0.32%) 0.0415 (+0.97%) 3.0547 (+3.08%) 4.7966 (-5.31%) 9.4243 (+2.29%)
3DGS [23] Wav2vec 2.0 [3] 30.8003 (-0.42%) 0.0421 (+2.43%) 3.2513 (+9.72%) 3.9522 (-21.9%) 9.6083 (+4.29%)
Whisper [15] 30.5449 (-0.28%) 0.0410 (-0.24%) 3.1397 (+5.93%) 3.2848 (-35.1%) 10.344 (+12.3%)
SE4Lip(Ours) 31.0176 (+0.29%) 0.0401 (-2.43%) 2.6836 (-9.44%) 8.3487 (+64.8%) 6.6705 (-27.6%)
Table 3: The quantitative results of video synthesis using different speech features. We highlight the best and second best results.
The extracted STFT spectrogram is processed through a Gated Recurrent Unit (GRU) network. The core of GRU consists of the update gate $z_{t}$ and the reset gate $r_{t}$. For the STFT spectrogram input sequence$\{x_{1},...,x_{T}\}$, the definitions of $z_{t}$ and $r_{t}$ are given in Eq. 1 and Eq. 2.
$z_{t}=\sigma(W_{z}\cdot[h_{t-1},x_{t}])$ (1)
$r_{t}=\sigma(W_{r}\cdot[h_{t-1},x_{t}])$ (2)
Here, $\sigma$ denotes the Sigmoid function, $h_{t-1}$ is the hidden state at the previous time step, and $W_{z}$ and $W_{r}$ are the corresponding weight matrices. The computation process of GRU at time step $t$ is represented by Eq. 3 and Eq. 4.
$\tilde{h}_{t}=\tanh(W_{h}\cdot[r_{t}\odot h_{t-1},x_{t}])$ (3)
$h_{t}=(1-z_{t})\odot h_{t-1}+z_{t}\odot\tilde{h}_{t}$ (4)
Here, $\tilde{h}_{t}$ is the candidate hidden state, and $h_{t}$ is the final hidden state. We select the output of the last GRU layer as the final feature. This network effectively models the temporal feature of speech. It captures the long-term dynamic changes in speech. This is particularly important in the talking head synthesis task, where the temporal relationships between phonemes are crucial. Therefore, we choose GRU instead of the commonly used CNN to better model these temporal dependencies. Through the output of the GRU, we obtain a speech embedding that contains dynamic information, providing strong support for the subsequent alignment of the rendering model.
4.3 Lip Feature Extractor
We aim to provide accurate lip features for aligning speech and lip shapes. Inspired by works [30, 7], we designed a CNN-based model to extract lip features. The specific architecture of this model is shown in Tab. 2. The model captures fine-grained spatial features of the lip images layer by layer through multiple convolutional layers. After processing the lip shape image through the CNN, an embedding is obtained with the same dimension as the speech features, representing the dynamic visual features of the lip shapes.
4.4 Loss Function
We use cosine similarity to measure the similarity between speech embedding and lip shape embedding as Eq. 5.
$cos(a,v)=\frac{a\cdot v}{||a||||v||}$ (5)
Here, $a$ and $v$ represent the speech and lip shape embeddings. For each speech-lip feature pair, we use cross-entropy loss as the contrastive loss function. The loss function encourages correctly paired speech and lip features to be closer to each other in the feature space, while incorrectly paired speech and lip features are pushed farther apart. The specific loss function is shown in Eq. 6.
$\mathcal{L}=-y\cdot\log(cos(a,v))-(1-y)\cdot\log(1-cos(a,v))$ (6)
Here, $y$ is the pairing label, indicating whether the speech and lip shape pair match. In the joint embedding space, the distance between matching speech-lip pairs decreases while the distance between mismatched pairs increases.
Figure 6: The qualitative results of video synthesis using different speech features. Our method has the most accurate lip shapes and achieves the best visual quality.
5 Experiments
5.1 Experimental Settings
Dataset. We use the same well-edited video sequences from [29, 23], which contains 6072 frames in total. The video has a frame rate of 25 fps, a resolution of 512x512, and a subject is centered in the video.
Comparison Method. We selected the latest representative works, SyncTalk [29] and TalkingGaussian [23], as rendering models. Specifically, SyncTalk refers to the NeRF-based framework, while TalkingGaussian refers to the 3DGS-based framework. We compare the video quality driven by DeepSpeech [1], HuBERT [15], Wav2Vec2.0 [3], Whisper [31] and SE4Lip.
Implementation Details. SE4Lip uses STFT with hyperparameters is set to n-fft=512, win-length=512, and hop-length=128. The number of GRU layers is set to 8. During training, the learning rate is 5e-5, the window is 5, and the batch size is 16. For SyncTalk, we use SmoothL1 Loss [10] as the loss function. The training steps are set to 60,000, with the fine-tuning steps set to 88,000. For TalkingGaussian, we use L1 Loss as the loss function, and the training steps are set to 20,000. Both training and video rendering are performed on a single NVIDIA RTX 4090 GPU.
Audio A Audio B
LMD↓ LSE-C↑ LSE-D↓ LMD↓ LSE-C↑ LSE-D↓
HuBERT [15] 3.0471 6.5291 7.9572 3.0328 6.8863 7.8993
DeepSpeech [1] 3.1169 6.7682 7.6824 3.1052 6.3795 8.0862
Wav2Vec 2.0 [3] 3.7019 3.8128 9.8629 3.6718 3.5824 10.1847
Whisper [31] 3.7177 4.0613 9.5127 3.6791 3.7153 10.9471
SE4Lip(Ours) 2.9304 8.1294 6.9714 2.9362 8.0932 7.0046
Table 4: The quantitative results of lip sync accuracy. We used two different speech samples to synthesize the videos. We highlight the best and second best results.
5.2 Quantitative Evaluation
Metrics. For image fidelity, we use the Peak Signal-to-Noise Ratio (PSNR) to measure overall quality and the Learned Perceptual Image Patch Similarity (LPIPS) [44] to measure fine details. Additionally, we evaluate the accuracy of lip shapes using the landmark distance (LMD), Lip Sync Error Confidence (LSE-C) [30], and Lip Sync Error Distance (LSE-D) [30].
Evaluation Results. We present the quantitative results in Tab. 3. Our method exhibits significant advantages in lip sync accuracy. Our method’s LMD surpasses the four acoustics features. In the NeRF model, our method improves LSE-C and LSE-D by 13.7% and 14.2%, respectively, compared to the best-performing baseline (HuBERT [15]). This advantage is further amplified in the 3DGS model. Notably, the LSE-C and LSE-D of our method are very close to the ground truth video, indicating the fine-grained capture of speech dynamic information by the STFT-GRU combination and the effectiveness of the cross-modal alignment framework. Our method also achieves the best performance in the visual quality metrics. Additionally, we used out-of-distribution speech to drive video synthesis in the NeRF model. We present the experimental results in Tab. 4. We used LMD, LSE-C, and LSE-D as metrics. The experimental results show that our method outperforms the four acoustics features in terms of lip sync accuracy.
Method DeepSpeech [1] HuBERT [15] Wav2Vec 2.0 [3] Whisper [31] SE4Lip(Ours)
Lip-sync Accuracy 2.89 3.17 2.54 2.49 4.27
Image Quality 3.82 3.78 3.89 3.69 4.02
Video Realness 3.38 3.42 3.37 3.19 3.82
Table 5: The results of user study. The rating is on a scale of 1-5. A higher scores indicate better performance. We highlight the best and second best results.
5.3 Qualitative Evaluation
Evaluation Results. To intuitively assess the quality of the synthesized video, we present a comparison between our method and other speech features in Fig. 6. As shown, SE4Lip presents more accurate lip shapes and higher visual quality. In the right half of Fig. 6, we provide a detailed comparison between DeepSpeech [1] and our method. When the movement amplitude of the lip is large, DeepSpeech generates a blurry lip shape (as shown in the first row) and also synthesizes inaccurate lip shape (as shown in the second row). When the lips are fully closed, DeepSpeech reveals the teeth and fails to fully close the lips (as shown in the third row). Our method can closely approach the ground truth video. This is due to the cross-modal alignment framework, which effectively addresses the phoneme-viseme alignment ambiguity issue. And fine-grained speech feature modeling preserves richer feature details. Furthermore, we observe that inaccurate lip features also affect other parts of the image. For example, in the second row of the right half of the figure, the neck area of the subject shows a noticeable shadow when using DeepSpeech.
In addition, we also present the synthesis results of phonemes /d/ and /t/ in Fig. 7. As shown, due to the phoneme-viseme alignment ambiguity issue, HuBERT [15] does not align /d/ and /t/ to a similar lip shape. Conversely, due to the design of cross-modal alignment, our approach effectively addresses this issue.
User Study. We conducted a user study to assess the quality of the synthesized videos effectively. We sampled 20 video clips from the quantitative evaluation and invited 12 volunteers to participate in the study. We used the mean opinion score (MOS) as the metric. Volunteers were asked to rate the synthesized videos on three aspects: 1) Lip-sync Accuracy, 2) Image Quality, and 3) Video Realness. The average scores for each method are presented in Tab. 5. As shown, our method significantly outperforms the comparison methods in terms of Lip-sync Accuracy, indicating the effectiveness of the cross-modal alignment framework.
Figure 7: The qualitative results of phonemes /d/ and /t/. HuBERT [15] exhibits the phoneme-viseme alignment ambiguity issue, while our approach avoids this issue.
5.4 Ablation Study
NeRF 3DGS
PSNR↑ LPIPS↓ LMD↓ PSNR↑ LPIPS↓ LMD↓
STFT+GRU(Ours) 32.2302 0.0399 2.8725 31.0176 0.0401 2.6836
Mel+GRU 31.7962 0.0419 3.1194 30.4452 0.0425 2.8709
Mel+CNN 31.7617 0.0417 3.0929 30.6923 0.0402 2.9109
STFT+CNN 31.8757 0.0413 3.0089 30.7702 0.0400 2.7432
Table 6: The results of ablation study on speech spectrogram and modeling. We highlight the best and second best results.
To evaluate the effectiveness of the SE4Lip framework, we conducted ablation experiments from two dimensions: speech spectrograms and modeling. The experimental results are presented in Tab. 6. As shown, the combination of STFT spectrogram and GRU significantly outperforms other variants in both video fidelity and lip sync accuracy, highlighting the necessity of the collaborative design of modules in SE4Lip.
[…]
Modeling Comparison. The temporal characteristics of GRU offer an improvement over the static CNN encoder. Under the 3DGS rendering model, the LMD of STFT+GRU is reduced by 2.2% compared to STFT+CNN. The update gate mechanism of GRU can adaptively adjust the granularity of temporal modeling, while CNN’s fixed receptive field fails to capture dynamic changes at the phoneme level. The experiments show that joint optimization of temporal modeling and full-band spectrograms is crucial for improving lip sync accuracy.
5.5 Disscussion
Our experimental results have demonstrated that SE4Lip effectively improves the quality of talking head synthesis, particularly in terms of lips. Although we have implemented SE4Lip using relatively simple models, its effectiveness highlights that the phoneme-viseme alignment ambiguity we discovered is a significant barrier to synthesizing high-quality lips. Additionally, phoneme-viseme alignment ambiguity in other languages also deserves attention.
6 Conclusion
To address the issue of phoneme-viseme alignment ambiguity in talking head synthesis tasks, we propose the SE4Lip. SE4Lip aligns the speech with the lip shape through a cross-modal alignment framework rather than aligning linguistic representations as in acoustics features. Additionally, SE4Lip processes the speech using a combination of an STFT spectrogram and a GRU-based model. This approach effectively captures fine-grained features in the time and frequency domain, providing strong support for subsequent rendering models. Experimental results show that SE4Lip achieves state-of-the-art performance on both NeRF and 3DGS rendering models. Notably, in terms of lip sync accuracy, SE4Lip improves LSE-C and LSE-D by 13.7% and 14.2%, compared to the best baseline, and produces results close to the ground truth videos. Ablation experiments further present the effectiveness of the STFT-GRU combination.